# Measuring the splashback feature: Dependence on halo properties and history

### v1

@@ Abstract
In this study, we define the novel splashback depth $\mathcal{D}$ and width $\mathcal{W}$ to examine how the splashback features of dark matter haloes are affected by the physical properties of haloes themselves. We use the largest simulation run in the hydrodynamic MillenniumTNG project. By stacking haloes in bins of halo mass, redshift, mass-dependent properties such as peak height and concentration, and halo formation history, we measure the shape of the logarithmic slope of the density profile of dark matter haloes. Our results show that the splashback depth has a strong dependence on the halo mass which follows a power law $\mathcal{D}\propto\left(\log_{10}M\right)^{2.8}$. Properties with strong correlation with halo mass demonstrate similar dependence. The splashback width has the strongest dependence on halo peak height and follows a power law $\mathcal{W}\propto\nu^{-0.87}$. We provide the fitting functions of the splashback depth and width in terms of halo mass, redshift, peak height, concentrations and halo formation time. The depth and width are therefore considered to be a long term memory tracker of haloes since they depend more on accumulative physical properties, e.g., halo mass, peak height and halo formation time. They are shaped primarily by the halo’s assembly history, which exerts a stronger influence on the inner density profile than short-term dynamical processes. In contrast, the splashback features have little dependence on the short term factors such as halo mass accretion rate and most recent major merger time. The splashback depth and width can therefore be used to complement information gained from quantities like the point of steepest slope or truncation radius to characterise the halo’s history and inner structure.
@@
thanks: E-mail: qiaorong.yu@balliol.ox.ac.uk
@@ 1 Introduction
In idealized spherical collapse models, the apocentre of all particles occurs at the same radius, creating a caustic in the density profile (Huss, Jain & Steinmetz, 1999; Adhikari, Dalal & Chamberlain, 2014). This manifests itself as a minimum of the density profile (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014), although there is not a one-to-one match in practice. This has led to many studies using the point of steepest slope, $R_{\rm st}$, as a more easily measured proxy for measuring $R_{\rm sp}$. However, the relationship between $R_{\rm sp}$ and $R_{\rm st}$ is not exact, and there remains some debate over the use of $R_{\rm st}$ as a measurement for $R_{\rm sp}$ (Diemer, 2020). Additionally, the steepest slope of a density profile does not necessarily indicate the presence of the splashback radius; any halo will have a steepest point in its density profile. The splashback feature is associated with the orbit of matter in a halo and where the logarithmic slope decreases significantly, typically below a value of 3-3- 3 as described in an NFW profile (Navarro, Frenk & White, 1997).
The feature has also been detected using galaxies as a tracer of dark matter haloes in clusters (e.g. More et al., 2016; Baxter et al., 2017; Chang et al., 2018; Shin et al., 2019; Murata et al., 2020). Because it is not possible to track the trajectories of galaxies in observations, these studies use $R_{\rm st}$ of the galaxy number density as a proxy for $R_{\rm sp}$. In addition, the galaxy profile does not perfectly align with the dark matter density profile espcially depending on the population of galaxies used (O’Neil et al., 2022). Observations also view clusters in projection. Although effects from projecting the profiles can be mitigated with proper fitting, there is a still a decrease in the value of $R_{\rm st}$ obtained from profiles in projection compared to 3D profiles (Sun et al., 2025).
@@
The depth and width of the splashback feature, describing the sharpness and extent of the density gradient at the halo edge, may provide a comprehensive characterization of halo boundaries. However, they remain largely unexamined in the context of their dependence on halo properties, growth and environmental influences. The evolution of these features with redshift could also illuminate changes in halo dynamics and the surrounding cosmic web. Zhang et al. (2023), for example, noted the splashback feature can impact cluster mass estimates using weak lensing.
@@
The rest of this paper is structured as follows. In Section 2, we describe the simulations, halo selection, the definition of the splashback depth and width, and methods for identifying splashback features. In Section 3.1, we examine the splashback features as a function of halo mass and redshift. We look further into the splashback features dependence on halo accretion rate and halo merger history in Section 3.2 and 14 including halo peak height (3.3.2), concentration (3.3.1) as well as formation time (3.3.3). We summarise our conclusions in Section 4.
@@ 2.1 Simulations
In this work, we use data from the MillenniumTNG (MTNG) project (Hernández-Aguayo et al., 2023; Kannan et al., 2023; Bose et al., 2023; Pakmor et al., 2023). MTNG builds on the Illustris (Vogelsberger et al., 2014a, b) and IllustrisTNG simulations (Marinacci et al., 2018; Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018; Naiman et al., 2018) and combines it with the much larger volume reached by the iconic dark matter-only Millennium simulation. The hydrodynamic simulations are performed using the moving-mesh code Arepo (Springel, 2010; Weinberger, Springel & Pakmor, 2020). We focus primarily on the hydrodynamic version in this work, and we briefly compare to the dark matter-only in Appendix B.
@@
Halo structures are identified using the friends-of-friends (FoF) algorithm (Davis et al., 1985), while subhaloes are detected with the SubFind algorithm (Springel et al., 2001; Dolag et al., 2009). The algorithm identifies structures by linking particles with neighbouring particles that lie within a linking length of $b=0.2$. The most massive gravitationally bound object in a FoF group is the main halo and others are labelled as subhaloes. The halo centre is defined as the position of the most bound particle within the main halo by using SubFind. To trace the formation histories of haloes, MTNG employs the SubLink merger tree algorithm, which links subhaloes across snapshots based on their most bound particle. This enables a consistent and physically motivated tracking of halo assembly over cosmic time (Rodriguez-Gomez et al., 2015).
@@ 2.2 Halo properties
We investigate the dependence of splashback features on the physical properties of dark matter haloes. This includes the halo mass, mass accretion rate, mass-dependent properties such as the peak height and concentration, as well as halo formation history. In this section, we clarify the definitions of the physical quantities that we use in the study.
@@ 2.2.2 Halo mass accretion rate
[…]
To compute the accretion rate of a halo at a given redshift, we first calculate the dynamical time at this redshift, then select the snapshot earlier in the simulation with a time difference closest to the calculated dynamical time. We note that we do not calculate the instantaneous accretion rate but the time-averaged accretion rate over the dynamical time to reduce noise from events like mergers. We use the SubLink merger trees (Rodriguez-Gomez et al., 2015) to identify the main progenitor of the halo. This algorithm tracks subhaloes through the simulation by linking subhaloes in consecutive snapshots that share the most bound particles. Therefore, by identifying the parent halo of the linked subhaloes, we can trace the history of haloes and then calculate the change in halo mass every dynamical time. By dividing the change in halo mass by the change in scale factor, we get the time-averaged halo accretion rate.
@@ 2.2.3 The most recent major merger
The most recent time of halo merging is important for shaping a halo’s dynamical state, internal structure, and density profile (Diemer, 2017). We stack haloes by their most recent major merging time and compute the median density profile. We define a major merger to be a merger where the smaller halo is at least 10 percent of the larger halo’s mass. The SubLink merger trees (Rodriguez-Gomez et al., 2015) is also used for this calculation. We cycle through the snapshots starting at $z=0$ and identify the main and second progenitors. When the ratio between the two exceeds 0.1, we record the snapshot where the two haloes merged as the time of the most recent major merger. Specifically, if the second and the main progenitor mass ratio exceeds 0.10.10.1 for the last time at the snapshot $n$, we choose snapshot $n+1$ as the time when merging occurs. It may be possible for a merger with a ratio of just under 0.10.10.1 to have been larger than this threshold in the past but shrank due to stripping of the smaller halo, but we maintain our method for simplicity.
@@ 2.2.4 Concentration
in which the scale radius $r_{s}$ is from the NFW profile (Navarro, Frenk & White, 1996):
@@
By fitting the density profile of individual halo with NFW profile within the range $r=0.01R_{\rm 200\text{m}}$ and $r=R_{\rm 200\text{m}}$, we calculate the optimal parameters $\rho_{\mathrm{s}}$ and $r_{\mathrm{s}}$, and then compute the concentration of each halo. Haloes are therefore stacked by concentration and the median density profile of every concentration stack can be computed.
@@ 2.2.5 Peak height
The peak height of a halo relates to how overdense a halo is and has direct correlation with halo mass and redshift (Press & Schechter, 1974). We use Colossus (Diemer, 2018) to calculate the peak height for a halo with a given mass and redshift. The peak height $\nu$ is defined as
@@
where $\delta_{c}\simeq 1.686$ is the linear overdensity threshold for halo collapse and $\sigma(M,z)$ is the rms variance of the linear density field on a scale for mass $M$ at redshift $z$.
@@ 2.2.6 Formation time
There are various ways of examining the formation time of haloes, including definitions based on the redshift when the halo reaches a certain fraction of its peak mass (Wechsler et al., 2002), or the time when its main progenitor first exceeds a fixed mass threshold (Giocoli et al., 2010). We define the halo formation time as the redshift when halo reaches half of its present mass $M_{200\text{m}}$. It is the snapshot at which the mass ratio $M_{200{\text{m}}}(z)$ and $M_{200{\text{m}}}(0)$ is closest to 0.5.
@@ 2.3 Halo samples
We also bin haloes by other features, e.g. accretion rate. We use the same halo samples summarised in Table 1. Instead of binning by mass, we calculate the accretion rate of each halo at each redshift using Equation 2 and bin haloes by accretion rate. The same strategy is applied to any other physical quantities we investigate.
Redshift Mass bins $[{\rm M}_{\odot}]$ Total
@@ 2.4 Fitting
To identify the splashback features of dark matter haloes, we stack haloes and take the median density as a function of radius to obtain a clean density profile. We identify the splashback feature as the region around the minimum of the logarithmic slope and bootstrap to obtain errors. We define the depth the feature as the height of the feature from the minimum of the logarithmic slope to where the slope flattens at larger radius and the width of the feature at the midpoint of the depth. We use an empirical function that models both the inner and outer halo regions of the density profile. We then compute the logarithmic slope of the fitted profiles and define the splashback features. This section describes the fitting procedure, and the next section gives details about how we estimate the splashback features from the fitted logarithmic slope of the fitted profiles.
@@ 2.4.1 The density profiles
Individual haloes are not themselves spherically symmetric objects, and the averaged density profiles can be noisy, especially for low-mass haloes and at the outskirts of a halo. Therefore, we stack sets of haloes depending on the property being examined (e.g. mass or accretion rate). When grouping by mass, we separate the haloes into five logarithmically spaced mass bins, with $\log_{10}(M_{200\text{m}}/{M_{\odot}})$ between 13, 13.5, 14, 14.5. 15, and 15.5 respectively. For accretion rate, we select haloes with accretion rate $0\leq\Gamma\leq 6$ and separate the haloes into 12 accretion rate bins, with $\Gamma$ between 0, 0.5, 1, 1.5, …, 4.5, 5, 5.5, and 6. Before stacking haloes, we normalise the profiles by $R_{\rm 200m}$ to mitigate effects from size variations smearing out the splashback feature. We take the median average density value in each radial bin to yield a stacked density profile as a function of median $R_{200\text{m}}$. The top panel of Figure 1 is an example density profiles of dark matter halo stacks with a mass of $10^{14}-10^{14.5}M_{\odot}$ at $z=0$.
@@ 2.4.2 Fitting the density profiles
Following O’Neil et al. (2021), we fit the slope directly by taking the numerical logarithmic derivative of the stacked halo profiles and the logarithmic derivative of Equation (7). We compute the numerical derivative of the density profiles using a central finite difference scheme with fourth-order accuracy for the logarithmic profile:
@@
We use the curve_fit function in the SciPy package to fit the slope profiles. We start with fitting the median density profile of every stack with the parameters described in Equation 7. We then use the found optimal parameters as the initial guess and fit the gradient profile with Equation 10. There has also been a simpler equation proposed to fit halo density profiles out to large radius in Diemer (2023). We explore how altering this fitting function affects our results in Appendix A.
@@ 2.4.3 Error estimation
These 10000 haloes samples are then separated into bins depending on what physical quantities we are looking at. For example, when binning the 10000 haloes by mass, we separate them into the five predefined mass bins. If there are no haloes in any bin whose original samples have haloes, we discard that sample and create another sample with 10000 haloes until all bins contain at least 1 halo. For example, at $z=0$, all five mass bins contain haloes, so we bootstrap 10,000 haloes across all bins. At $z=2$, only three bins are populated (see Table 1), so we ensure that the lowest three bins contain haloes with each bootstrap. This is counted as one effective sampling. In this case, we move on to calculate the splashback features of the stacked haloes in each bin. We repeat the effective sampling 1024 times. For each bin, there are therefore 1024 splashback features calculated. We assume a Gaussian prior to these values and choose the median value as the splashback feature and take the $16^{\text{th}}$ and $84^{\text{th}}$ percentiles, approximating a standard deviation of a Gaussian distribution, as the errors of the splashback feature.

### v2

@@ Abstract
In this study, we define the novel splashback depth $\mathcal{D}$ and width $\mathcal{W}$ to examine how the splashback features of dark matter haloes are affected by the physical properties of haloes themselves. We use the largest simulation run in the hydrodynamic MillenniumTNG project. By stacking haloes in bins of halo mass, redshift, mass-dependent properties such as peak height and concentration, and halo formation history, we measure the shape of the logarithmic slope of the density profile of dark matter haloes. Our results show that the splashback depth has a strong dependence on the halo mass which follows a power law $\mathcal{D}\propto\left(\log_{10}M\right)^{2.8}$. Properties with strong correlation with halo mass demonstrate similar dependence. The splashback width has the strongest dependence on halo peak height and follows a power law $\mathcal{W}\propto\nu^{-0.87}$. We provide the fitting functions of the splashback depth and width in terms of halo mass, redshift, peak height, concentrations and halo formation time. The depth and width are therefore considered to be a long-term memory tracker of haloes since they depend more on accumulative physical properties, e.g., halo mass, peak height and halo formation time. They are shaped primarily by the halo’s assembly history, which exerts a stronger influence on the inner density profile than short-term dynamical processes. In contrast, the splashback features have little dependence on the short-term factors such as halo mass accretion rate and most recent major merger time. The splashback depth and width can therefore be used to complement information gained from quantities like the point of steepest slope or truncation radius to characterise the halo’s history and inner structure.
@@
thanks: E-mail: shannon.qyu@gmail.com
@@ 1 Introduction
In idealised spherical collapse models, the density caustic arises from the pile-up of infalling shells near their first apocentre, where the apocentre of each particle’s first orbit occurs at roughly the same radius (Huss, Jain & Steinmetz, 1999; Adhikari, Dalal & Chamberlain, 2014). This manifests itself as a minimum of the density slope profile (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014), although there is not a one-to-one match in practice. This has led to many studies using the point of steepest slope, $R_{\rm st}$, as a more easily measured proxy for measuring $R_{\rm sp}$. However, the relationship between $R_{\rm sp}$ and $R_{\rm st}$ is not exact, and there remains some debate over the use of $R_{\rm st}$ as a measurement for $R_{\rm sp}$ (Diemer, 2020). Additionally, the steepest slope of a density profile does not necessarily indicate the presence of the splashback radius; any halo will have a steepest point in its density profile. The splashback feature is associated with the orbit of matter in a halo and where the logarithmic slope decreases significantly, typically below a value of 3-3 as described in an NFW profile (Navarro, Frenk & White, 1997).
The feature has also been detected using galaxies as a tracer of dark matter haloes in clusters (e.g. More et al., 2016; Baxter et al., 2017; Chang et al., 2018; Shin et al., 2019; Murata et al., 2020). Because it is not possible to track the trajectories of galaxies in observations, these studies use $R_{\rm st}$ of the galaxy number density as a proxy for $R_{\rm sp}$. In addition, the galaxy profile does not perfectly align with the dark matter density profile, especially depending on the population of galaxies used (O’Neil et al., 2022). Observations also view clusters in projection. Although effects from projecting the profiles can be mitigated with proper fitting, there is still a decrease in the value of $R_{\rm st}$ obtained from profiles in projection compared to 3D profiles (Sun et al., 2025).
@@
The depth and width of the splashback feature, describing the sharpness and extent of the density gradient at the halo edge, may provide a comprehensive characterisation of halo boundaries. However, they remain largely unexamined in the context of their dependence on halo properties, growth and environmental influences. The evolution of these features with redshift could also illuminate changes in halo dynamics and the surrounding cosmic web. Zhang et al. (2023), for example, noted the splashback feature can impact cluster mass estimates using weak lensing.
@@
The rest of this paper is structured as follows. In Section 2, we describe the simulations, halo selection, the definition of the splashback depth and width, and methods for identifying splashback features. In Section 3.1, we examine the splashback features as a function of halo mass and redshift. We look further into the splashback feature’s dependence on halo accretion rate and halo merger history in Sections 3.2 and 14 including halo peak height (3.3.2), concentration (3.3.1) as well as formation time (3.3.3). We summarise our conclusions in Section 4.
@@ 2.1 Simulations
In this work, we use data from the MillenniumTNG (MTNG) project (Hernández-Aguayo et al., 2023; Kannan et al., 2023; Bose et al., 2023; Pakmor et al., 2023). MTNG builds on the Illustris (Vogelsberger et al., 2014a, b) and IllustrisTNG simulations (Marinacci et al., 2018; Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018; Naiman et al., 2018) and combines them with the much larger volume reached by the iconic dark matter-only Millennium simulation. The hydrodynamic simulations are performed using the moving-mesh code Arepo (Springel, 2010; Weinberger, Springel & Pakmor, 2020). We focus primarily on the hydrodynamic version in this work, and we briefly compare it to the dark matter-only version in Appendix B.
@@
Halo structures are identified using the friends-of-friends (FoF) algorithm (Davis et al., 1985), while subhaloes are detected with the SubFind algorithm (Springel et al., 2001; Dolag et al., 2009). The algorithm identifies structures by linking particles with neighbouring particles that lie within a linking length of $b=0.2$. The most massive gravitationally bound object in an FoF group is the main halo and others are labelled as subhaloes. The halo centre is defined as the position of the most bound particle within the main halo by using SubFind. To trace the formation histories of haloes, MTNG employs the SubLink merger tree algorithm, which links subhaloes across snapshots based on their most bound particle. This enables a consistent and physically motivated tracking of halo assembly over cosmic time (Rodriguez-Gomez et al., 2015).
@@ 2.2 Halo properties
We investigate the dependence of splashback features on the physical properties of dark matter haloes. This includes the halo mass, mass accretion rate, and mass-dependent properties such as the peak height and concentration, as well as halo formation history. In this section, we clarify the definitions of the physical quantities that we use in the study.
@@ 2.2.2 Halo mass accretion rate
where $t_{H}$ is the Hubble time $1/H(z)$. To compute the accretion rate of a halo at a given redshift, we first calculate the dynamical time at this redshift, then select the snapshot earlier in the simulation with a time difference closest to the calculated dynamical time. We note that we do not calculate the instantaneous accretion rate but the time-averaged accretion rate over the dynamical time to reduce noise from events like mergers. We use the SubLink merger trees (Rodriguez-Gomez et al., 2015) to identify the main progenitor of the halo. This algorithm tracks subhaloes through the simulation by linking subhaloes in consecutive snapshots that share the most bound particles. Therefore, by identifying the parent halo of the linked subhaloes, we can trace the history of haloes and then calculate the change in halo mass every dynamical time. By dividing the change in halo mass by the change in scale factor, we get the time-averaged halo accretion rate.
@@ 2.2.3 The most recent major merger
The most recent time of halo merging is important for shaping a halo’s dynamical state, internal structure, and density profile (Diemer, 2017). We stack haloes by their most recent major merging time and compute the median density profile. We define a major merger to be a merger where the smaller halo is at least 10 percent of the larger halo’s mass. The SubLink merger trees (Rodriguez-Gomez et al., 2015) are also used for this calculation. We cycle through the snapshots starting at $z=0$ and identify the main and second progenitors. When the ratio between the two exceeds 0.1, we record the snapshot where the two haloes merged as the time of the most recent major merger. Specifically, if the second and the main progenitor’s mass ratio exceeds $0.1$ for the last time at the snapshot $n$, we choose the snapshot $n+1$ as the time when merging occurs. It may be possible for a merger with a ratio of just under $0.1$ to have been larger than this threshold in the past but shrunk due to stripping of the smaller halo, but we maintain our method for simplicity.
@@ 2.2.4 Concentration
in which the scale radius $r_{s}$ is from the NFW profile (Navarro, Frenk & White, 1996)
@@
By fitting the density profile of individual halo with the NFW profile within the range $r=0.01R_{\rm 200\text{m}}$ and $r=R_{\rm 200\text{m}}$, we calculate the optimal parameters $\rho_{\mathrm{s}}$ and $r_{\mathrm{s}}$, and then compute the concentration of each halo. Haloes are therefore stacked by concentration and the median density profile of every concentration stack can be computed.
@@ 2.2.5 Peak height
The peak height of a halo relates to how overdense a halo is and has a direct correlation with halo mass and redshift (Press & Schechter, 1974). The peak height $\nu$ provides a dimensionless measure of the rarity of the initial density peak from which a halo forms. It compares the critical overdensity for spherical collapse $\delta_{c}$, with the RMS variance $\sigma\left(M,z\right)$ of the linear density field smoothed on the mass scale $M$. Halos with the same peak height $\nu$ arise from equally rare peaks in the primordial density field. They are therefore at comparable stages of gravitational collapse and reside in statistically similar environments, independent of redshift (e.g. Sheth & Tormen, 1999; Diemer & Kravtsov, 2015). Studies have shown that the splashback radius has an obvious correlation with peak height, so peak height can potentially correlate with other splashback features which clears out the redshift dependence (More, Diemer & Kravtsov, 2015; Diemer et al., 2017). We use Colossus (Diemer, 2018) to calculate the peak height for a halo with a given mass and redshift. The peak height $\nu$ is defined as
@@
where $\delta_{c}\simeq 1.686$ is the linear overdensity threshold for halo collapse, and $\sigma(M,z)$ is the RMS variance of the linear density field on a scale for mass $M$ at redshift $z$.
@@ 2.2.6 Formation time
There are various ways of examining the formation time of haloes, including definitions based on the redshift when the halo reaches a certain fraction of its peak mass (Wechsler et al., 2002), or the time when its main progenitor first exceeds a fixed mass threshold (Giocoli et al., 2010). We define the halo formation time as the redshift when the halo reaches half of its present mass $M_{200\text{m}}$. It is the snapshot at which the mass ratio $M_{200{\text{m}}}(z)$ and $M_{200{\text{m}}}(0)$ is closest to 0.5.
@@ 2.3 Halo samples
We also bin haloes by other features, e.g., accretion rate. We use the same halo samples summarised in Table 1. Instead of binning by mass, we calculate the accretion rate of each halo at each redshift using Equation 2 and bin haloes by accretion rate. The same strategy is applied to any other physical quantities we investigate.
Redshift Mass bins $\rm M$$_{\odot}$] Total
@@ 2.4 Fitting
To identify the splashback features of dark matter haloes, we stack haloes and take the median density as a function of radius to obtain a clean density profile. We identify the splashback feature as the region around the minimum of the logarithmic slope and bootstrap to obtain errors. We define the depth of the feature as the height of the feature from the minimum of the logarithmic slope to where the slope flattens at a larger radius and the width of the feature at the midpoint of the depth. We use an empirical function that models both the inner and outer halo regions of the density profile. We then compute the logarithmic slope of the fitted profiles and define the splashback features. This section describes the fitting procedure, and the next section gives details about how we estimate the splashback features from the fitted logarithmic slope of the fitted profiles.
@@ 2.4.1 The density profiles
Individual haloes are not themselves spherically symmetric objects, and the averaged density profiles can be noisy, especially for low-mass haloes and at the outskirts of a halo. Therefore, we stack sets of haloes depending on the property being examined (e.g., mass or accretion rate). When grouping by mass, we separate the haloes into five logarithmically spaced mass bins, with $\log_{10}(M_{200\text{m}}/{M_{\odot}})$ between 13, 13.5, 14, 14.5, 15, and 15.5, respectively. For accretion rate, we select haloes with accretion rates $0\leq\Gamma\leq 6$ and separate the haloes into 12 accretion rate bins, with $\Gamma$ between 0, 0.5, 1, 1.5, …, 4.5, 5, 5.5, and 6. Before stacking haloes, we normalise the profiles by $R_{\rm 200m}$ to mitigate effects from size variations smearing out the splashback feature. We take the median average density value in each radial bin to yield a stacked density profile as a function of median $R_{200\text{m}}$. The top panel of Figure 1 is an example of density profiles of dark matter halo stacks with a mass of $10^{14}-10^{14.5}M_{\odot}$ at $z=0$.
@@ 2.4.2 Fitting the density profiles
Following O’Neil et al. (2021), we fit the slope directly by taking the numerical logarithmic derivative of the stacked halo profiles and the logarithmic derivative of Equation (7). We compute the numerical derivative of the density profiles using a central finite difference scheme with fourth-order accuracy for the logarithmic profile
@@
We use the curve_fit function in the SciPy package to fit the slope profiles. We start with fitting the median density profile of every stack with the parameters described in Equation 7. We then use the found optimal parameters as the initial guess and fit the gradient profile with Equation 10. There has also been a simpler equation proposed to fit halo density profiles out to a large radius in Diemer (2023). We explore how altering this fitting function affects our results in Appendix A.
@@ 2.4.3 Error estimation
These 10000 halo samples are then separated into bins depending on what physical quantities we are looking at. For example, when binning the 10000 haloes by mass, we separate them into the five predefined mass bins. If there are no haloes in any bin whose original samples have haloes, we discard that sample and create another sample with 10000 haloes until all bins contain at least 1 halo. For example, at $z=0$, all five mass bins contain haloes, so we bootstrap 10,000 haloes across all bins. At $z=2$, only three bins are populated (see Table 1), so we ensure that the lowest three bins contain haloes with each bootstrap. This is counted as one effective sampling. In this case, we move on to calculate the splashback features of the stacked haloes in each bin. We repeat the effective sampling 1024 times. For each bin, there are therefore 1024 splashback features calculated. We assume a Gaussian prior to these values and choose the median value as the splashback feature and take the $16^{\text{th}}$ and $84^{\text{th}}$ percentiles, approximating a standard deviation of a Gaussian distribution, as the errors of the splashback feature.